OpenStudy (anonymous):

What is the least common denominator for the fractions 1/6 and 3/4 ? A. 18 B. 24 C. 12 D. 8

OpenStudy (anonymous):

so 6 = 100 and 4 = 100??

OpenStudy (anonymous):

@Liv1234

OpenStudy (anonymous):

1/6 of 100 = 16.6666666667

OpenStudy (liv1234):

Do you know what a common denominator is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3/4 of 100 = 33.3333333333

OpenStudy (liv1234):

Okay, so I know what the answer is, but I don't want to give you direct answers.

OpenStudy (anonymous):

im leaving this to you @Liv1234

OpenStudy (liv1234):

Okay, thank you.

OpenStudy (liv1234):

So, we know that 8 is not a multiple of 6, but it is a multiple of 4, which means that D, is not the answer.

OpenStudy (liv1234):

If we have 24, 18, and 12 left, which number is the lowest of the three?

OpenStudy (anonymous):

12

OpenStudy (liv1234):

Correct, so do you know what this means?

OpenStudy (anonymous):

the answer is 12

OpenStudy (liv1234):

So, C. And do you know why C is the answer?

OpenStudy (anonymous):

yes because 6*2=12 4*3=12

OpenStudy (liv1234):

Correct! If you need any other help with any other questions, open a new post and tag me.